OpenStudy (anonymous):

1 1/3 times 5 5/8 = ? 8 5 6/11 7 1/2 5 5/24

OpenStudy (anonymous):

make improper fractions out of both and cancel and multiply: 4/3 x 45/8=15/2=7 1/2